Sin inverse x +sin inverse y =pie then valueof square of (x+y)

Grade:12

1 Answers

Arun
6 years ago
We Know that
Sin-1x + cos-1x = \Pi/2
And Given is that Sin-1x + Sin-1y = \Pi/2
Therefore , we get that...
Sin-1y = cos-1x
Sin-1y = Sin-1(\sqrt{1 - x^{2}}
 
Therefore
y = \sqrt{1 - x^{2}}
 
On Squaring both sides .
we get
 
x?2  = 1 – y?2